# Analysing Zero-Shot Readability-Controlled Sentence Simplification

### v1

@@ 1 Introduction
Most methods for sentence-level RCTS rely on fully supervised training approaches (Nishihara et al., 2019; Spring et al., 2021; Rios et al., 2021; Yanamoto et al., 2022; Zetsu et al., 2022; Agrawal and Carpuat, 2023), and are hindered by the scarcity of high-quality annotated datasets that include original-simplified sentence pairs alongside readability levels. Parallel datasets such as Newsela (Xu et al., 2015) and OneStopEnglish (Vajjala and Lučić, 2018) contain readability annotations only at the document level, which cannot be directly transferred to the sentences in each text. To address this challenge, we use instruction-tuned large language models (LLMs) for RCTS on English sentences without parallel annotated data (i.e. in a zero-shot setting). We exploit LLMs’ ability to follow natural language instructions by providing different types of contextual information about the desired readability, including descriptions of readers’ skills at the target level and examples of sentences with the required readability.
We add to existing work on LLMs for RCST (Imperial and Tayyar Madabushi, 2023; Farajidizaji et al., 2023; Agrawal and Carpuat, 2023) by: (1) analysing models’ performance at the sentence level rather than the passage level; (2) examining a higher number of LLMs (with a total of 8 models), and (3) measuring not just a model’s ability to meet the target readability, but also how this affects preserving the original meaning using both automatic metrics and human assessment. We find that:
@@ 2.3 Instruction-Tuned LLMs
We investigate 8 instruction-tuned LLMs with diverse sizes and two types of architecture: an encoder-decoder and a standalone decoder, including open-weight and closed-weight models. For the open-weight models, we use LLMs in the 80M-8B range: Flan-T5 (Chung et al., 2022) in four size variants Small (80M), Base (250M), Large (780M), and XL (3B); Llama3-instruct-8B (AI@Meta, 2024); OpenChat_3.5-7B (Wang et al., 2023); and Mistral-7B-Instruct (Jiang et al., 2023). For closed models, we use GPT-3.5-Turbo and GPT-4-Turbo from OpenAI. [^4] https://platform.openai.com/docs/models More information on the models can be found in Appendix A.
@@ 3.3 Baselines
Table 2 presents the scores obtained by all selected models in our test set, with the different designed prompts, and across all automatic metrics. The following sections analyse the results in terms of achievement of target readability levels (Sec. 4) and preservation of the original meaning (Sec. 5).
@@ 4.2 The Effect of Source vs Target Readability
Figure 2 presents the RMSE scores for the top five performing models across all combinations of source and target CEFR levels. The heatmaps indicate that simplification becomes increasingly difficult (evidenced by higher RMSE) as target levels decrease. Models achieve their best performance when source and target CEFR levels are close, especially in simplifications from B2 to B1, as shown by the blue sections. This suggests that models handle moderate simplifications with fewer complexity adjustments more effectively. Conversely, models struggle most as the gap between source and target levels widens, especially when simplifying from C2 to A1, highlighting their limited ability to handle large reductions in complexity.
@@ 4.4 Do Humans Agree with the Metrics?
Table 3: Human readability assessment results for GPT-4 and LlaMA 3. AccuracyH: Human-rated accuracy; AccuracyA: Classifier-rated accuracy.
We randomly sampled 45 simplifications (15 per target level) from GPT-4 and Llama 3, totaling 90 instances. We recruited three evaluators with experience in teaching English as a second language, and asked them to: (1) assess the CEFR level that a student would need to understand each generated simplification, and (2) judge whether the corresponding source sentences could actually be adjusted to match the target CEFR levels. This aimed to establish if the fact that an output did not reach the target level was due to issues with the model itself or the nature of the source text. [^8] Details on evaluator selection, inter-rater agreement, and the exact instructions provided can be found in Appendix H.1.
@@
Table 3 shows that both GPT-4-Turbo and Llama 3-Instruct produced sentences that aligned with the target CEFR levels in 80% and 84.4% of cases, respectively, based on human evaluations (AccuracyH). This indicates that in the majority of cases, human raters judged the generated sentences to match closely the expected readability levels for the intended audience.
@@ 5.1 Vector-based Semantic Similarity
Through manual inspection, we found that STS tends to over-penalize text changes that could be considered appropriate in RCTS. In order to make a text more accessible, simplification can sometimes involve removing or inserting new information, such when executing compression or elaboration operations, respectively (Alva-Manchego et al., 2020). Such forms of rewriting might result in STS scores that underestimate the true quality of the simplifications. For example, the first three simplifications in Table 5 could be considered of high quality since they accurately capture the meaning of their original counterparts. However, they did not receive the nearly-perfect STS scores that the last conservative simplification did. Experiments with BERTscore (Zhang et al., 2020) and AlignScore (Zha et al., 2023), typically used for meaning preservation and factuality evaluation, yielded similar results: BERTscore favored outputs that minimally modify or directly copy the input text; while AlignScore penalized the addition of supporting details (such as elaboration and examples) deeming them incorrect, while rewarding factual accuracy even when simplifications are inappropriate. [^9] More details, examples, and results for BERTscore and AlignScore are available in Appendix G This highlights the necessity for meaning preservation metrics that account for the nuances of the editing operations executed while simplifying texts.
@@ 5.2 Human Judgements
We randomly sampled 45 simplifications from GPT-4 and Llama 3, totaling 90 instances. Three fluent English speakers evaluated whether the original meaning was retained in the simplified sentences. We adapted the error annotation scheme for evaluating factuality in TS proposed by Devaraj et al. (2022): instead of scoring each operation independently, we asked annotators to provide an overall score for each simplification. Annotators rated each simplification using the following scale: 0 indicates no or trivial changes; 1 signifies that non-trivial changes have occurred, but the main idea is maintained; and 2 indicates that the main idea has not been preserved in the simplified sentence. We also allowed intermediate scores (0.5 and 1.5). [^10] Details on evaluator selection, inter-rater agreement, and the instructions provided can be found in Appendix H.2.
@@ 7 Conclusion
We explored sentence-level zero-shot RCTS to analyse various instruction-tuned LLMs in simplifying sentences to specified CEFR levels. Our experiments showed that LLMs have difficulties with this task, especially when there was a large gap between source and target readability levels. Providing contextual information about the target readability (i.e. CEFR descriptions and example sentences) improves performance but not uniformly.
We also demonstrated that automatically assessing models for RCTS is challenging due to the nature of the task. Standard similarity metrics penalise removing or adding information even if it improves readability, while CEFR predictors often misjudge the true difficulty of simplified texts, failing to account for rewritings via elaboration and sentence splitting. This gap between human and automated assessments underscores the need for more nuanced evaluation mechanisms in RCTS.
@@
For readability assessments, evaluators were paid at \$65.99/h, a rate significantly higher than the local minimum wage, demonstrating our respect for the specialized expertise required. For meaning preservation assessments, evaluators received a fair compensation of \$33/h.
@@ Appendix E CEFR Level Assessment Performance
Table 9 provides a comparative analysis of CEFR level assessment performance on the publicly available CEFR-SP test set from SCoRE and Wiki-Auto. This comparison compares various LLMs including the GPT-4 model, OpenChat-3.5, and Llama-3-8B-Instruct, as well as the README sentence-level CEFR level estimator (Naous et al., 2023), and the CEFR sentence-based estimator used in this study Arase et al. (2022). The results indicate that the performance of these models was lower than that achieved by the model developed by Arase et al. (2022). Moreover, Tables 11, 12 and 13 show the prompt provided to the GPT-4, Llama-3-8B-Instruct and OpenChat-3.5 models for the assessment, respectively.
@@ Appendix G Meaning Preservation Supplemental Results
Table 15 presents the BERTScore and AlignScore results obtained by all selected models in our test set, with the different designed prompts. Moreover, Table 10 provides examples that illustrate the impact of simplification on AlignScore scores.

### v2

@@ 1 Introduction
Most methods for sentence-level RCTS rely on fully supervised training approaches (Nishihara et al., 2019; Spring et al., 2021; Rios et al., 2021; Yanamoto et al., 2022; Zetsu et al., 2022; Agrawal and Carpuat, 2023), and are hindered by the scarcity of high-quality annotated datasets that include original-simplified sentence pairs alongside readability levels. Parallel datasets such as Newsela (Xu et al., 2015) and OneStopEnglish (Vajjala and Lučić, 2018) contain readability annotations only at the document level, which cannot be directly transferred to individual paragraphs or sentences in each text. To address this, we investigate using instruction-tuned large language models (LLMs) for RCTS on English sentences without parallel annotated data (i.e. in a zero-shot setting). In particular, we use the CEFR-SP dataset  (Arase et al., 2022), since the CEFR standard offers multiple readability targets, which challenge controlled generation models. We exploit LLMs’ ability to follow natural language instructions by providing different types of contextual information about the desired readability, including descriptions of readers’ skills at the target CEFR level and examples of sentences with the required readability.
We add to existing work on LLMs for RCTS (Imperial and Tayyar Madabushi, 2023; Farajidizaji et al., 2023; Agrawal and Carpuat, 2023) by: (1) analysing models’ performance at the sentence level rather than the passage level; (2) examining a higher number of LLMs (with a total of 8 models), and (3) measuring not just a model’s ability to meet the target readability, but also how this affects preserving the original meaning using both automatic metrics and human assessment. We find that:
@@ 2.3 Instruction-Tuned LLMs
We investigate 8 instruction-tuned LLMs with diverse sizes and two types of architecture: an encoder-decoder and a standalone decoder, including open-weight and closed-weight models. For the open-weight models, we use LLMs in the 80M-8B range: Flan-T5 (Chung et al., 2022) in four size variants Small (80M), Base (250M), Large (780M), and XL (3B); Llama-3-instruct-8B (AI@Meta, 2024); OpenChat_3.5-7B (Wang et al., 2023); and Mistral-7B-Instruct (Jiang et al., 2023). For closed models, we use GPT-3.5-Turbo and GPT-4-Turbo from OpenAI. [^4] https://platform.openai.com/docs/models More information on the models can be found in Appendix A.
@@ 3.3 Baselines
Table 2 presents the scores obtained by all selected models in our test set, with the different designed prompts, and across all automatic metrics. The following sections analyze these results in terms of two core aspects: (1) the achievement of target readability levels (Sec. 4), which ensures the simplified text meets intended accessibility goals, and (2) the preservation of the original meaning (Sec. 5), essential for maintaining fidelity to the source content. This dual focus is crucial for evaluating the effectiveness of the models in producing accessible, yet accurate, simplified texts.
@@ 4.2 The Effect of Source vs Target Readability
Figure 2 presents the RMSE scores for the top five performing models across all combinations of source and target CEFR levels. The heatmaps indicate that simplification becomes increasingly difficult (evidenced by higher RMSE) as target levels decrease. Models achieve their best performance when source and target CEFR levels are close, especially in simplifications from B2 to B1, as shown by the blue sections. This suggests that models handle moderate simplifications with fewer complexity adjustments more effectively.
Conversely, models struggle most as the gap between source and target levels widens, especially when simplifying from C2 to A1, highlighting their limited ability to handle large reductions in complexity.
@@ 4.4 Do Humans Agree with the Metrics?
Table 3: Human readability assessment results for GPT-4-Turbo and Llama-3-8B-Instruct. AccuracyH: Human-rated accuracy; AccuracyA: Classifier-rated accuracy.
We randomly sampled 45 simplifications (15 per target level) from GPT-4-Turbo and Llama-3-8B-Instruct, totaling 90 instances. We recruited three evaluators with experience in teaching English as a second language, and asked them to: (1) assess the CEFR level that a student would need to understand each generated simplification, and (2) judge whether the corresponding source sentences could actually be adjusted to match the target CEFR levels. This aimed to establish if the fact that an output did not reach the target level was due to issues with the model itself or the nature of the source text. [^8] Details on evaluator selection, inter-rater agreement, and the exact instructions provided can be found in Appendix H.1.
@@
Table 3 shows that both GPT-4-Turbo and Llama-3-8B-Instruct produced sentences that aligned with the target CEFR levels in 80% and 84.4% of cases, respectively, based on human evaluations (AccuracyH). This indicates that in the majority of cases, human raters judged the generated sentences to match closely the expected readability levels for the intended audience.
@@ 5.1 Vector-based Semantic Similarity
Through manual inspection, we found that STS tends to over-penalize text changes that could be considered appropriate in RCTS. In order to make a text more accessible, simplification can sometimes involve removing or inserting new information, such when executing compression or elaboration operations, respectively (Alva-Manchego et al., 2020). Such forms of rewriting might result in STS scores that underestimate the true quality of the simplifications. For example, the first three simplifications in Table 5 could be considered of high quality since they accurately capture the meaning of their original counterparts. However, they did not receive the nearly-perfect STS scores that the last conservative simplification did. Experiments with BERTscore (Zhang et al., 2020) and AlignScore (Zha et al., 2023), typically used for meaning preservation and factuality evaluation, yielded similar results. As shown in Table 6, BERTscore favored outputs that minimally modify or directly copy the input text. On the other hand, AlignScore penalized the addition of supporting details (such as elaboration and examples) deeming them incorrect (as illustrated in the first and second examples in Table 7), while rewarding factual accuracy even when simplifications are inappropriate (as seen in the third example in Table 7). [^9] More details and results for BERTscore and AlignScore are available in Appendix G This highlights the necessity for meaning preservation metrics that account for the nuances of the editing operations executed while simplifying texts.
Original Simplified BERTscore
(C2) Pharmacoepidemiology bridges pharmacology and epidemiology and borrows their methodology from both of them. (B1) (GPT-4) Pharmacoepidemiology combines the study of drugs (pharmacology) and the study of health in populations (epidemiology). It uses methods from both these fields. 0.52
(C1) Rivlin had pulmonary fibrosis, a lung disease. (B1) (GPT-4) Rivlin had a lung disease called pulmonary fibrosis. 0.86
(C2) Dimethylmercury is extremely toxic and dangerous to handle. (B1) (Mistral-7B) Dimethylmercury is very harmful and should be handled with care. 0.75
(C2) Dimethylmercury is extremely toxic and dangerous to handle. (B1) (Flan-T5-L) Dimethylmercury is extremely dangerous and toxic to handle. 0.95
Table 6: Examples of the impact of simplification on BERTscore scores.
Original Simplified AlignScore
@@ 5.2 Human Judgements
We randomly sampled 45 simplifications from GPT-4-Turbo and Llama-3-8B-Instruct, totaling 90 instances. Three fluent English speakers evaluated whether the original meaning was retained in the simplified sentences. We adapted the error annotation scheme for evaluating factuality in TS proposed by Devaraj et al. (2022): instead of scoring each operation independently, we asked annotators to provide an overall score for each simplification. Annotators rated each simplification using the following scale: 0 indicates no or trivial changes; 1 signifies that non-trivial changes have occurred, but the main idea is maintained; and 2 indicates that the main idea has not been preserved in the simplified sentence. We also allowed intermediate scores (0.5 and 1.5). [^10] Details on evaluator selection, inter-rater agreement, and the instructions provided can be found in Appendix H.2.
@@ 7 Conclusion
We explored sentence-level zero-shot RCTS to analyse various instruction-tuned LLMs in simplifying sentences to specified CEFR levels. Our experiments showed that LLMs have difficulties with this task, especially when there is a large gap between source and target readability levels. Providing contextual information about the target readability (i.e. CEFR descriptions and example sentences) improves performance but not uniformly.
We also demonstrated that automatically assessing models for RCTS is challenging due to the nature of the task. Standard similarity metrics penalise removing or adding information even if it improves readability, while CEFR predictors often misjudge the true difficulty of simplified texts, failing to account for rewritings via elaboration and sentence splitting. This gap between human and automated assessments underscores the need for more nuanced evaluation mechanisms in RCTS. Refined metrics would not only enhance the evaluation process but also drive significant improvements in RCTS development by better supporting output filtering to ensure both readability and meaning preservation, or by serving as reward signals in reinforcement learning approaches.
@@
For readability assessments, we paid the evaluators \$65.99 for completing the task, which took an average of 1 hour to complete. Similarly, for meaning preservation assessments, evaluators received a flat rate of \$33 to complete the task, with an average task completion time of 40 minutes.
@@ Appendix E CEFR Level Assessment Performance
Table 11 provides a comparative analysis of CEFR level assessment performance on the publicly available CEFR-SP test set from SCoRE and Wiki-Auto. This comparison compares various LLMs including the GPT-4 model, OpenChat-3.5, and Llama-3-8B-Instruct, as well as the README sentence-level CEFR level estimator (Naous et al., 2023), and the CEFR sentence-based estimator used in this study Arase et al. (2022). The results indicate that the performance of these models was lower than that achieved by the model developed by Arase et al. (2022). Moreover, Tables 12, 13 and 14 show the prompt provided to the GPT-4, Llama-3-8B-Instruct and OpenChat-3.5 models for the assessment, respectively.
@@ Appendix G Meaning Preservation Supplemental Results
Table 16 presents the BERTScore and AlignScore results obtained by all selected models in our test set, with the different designed prompts.